# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ III Slow thermalization and the time complexity of the word problem
From the discussion of the previous section, it is natural to expect that systems with ${\sf Dyn}_{G}$ dynamics will have thermalization times controlled by the time complexity of the word problem, as diagnosed by the Dehn functions ${\sf Dehn}_{g}(L)$ defined in (14). In this section we make this relationship precise by using the functions ${\sf Dehn}_{g}(L)$ to place lower bounds on various thermalization timescales. In the case of random unitary or classical stochastic dynamics, ${\sf Dehn}_{g}(L)$ will be used to bound relaxation and mixing times; for Hamiltonian dynamics it will appear in bounds for hitting times. A common feature these timescales have is that they quantify when ${\sf Dyn}_{G}$ is able to spread initial product states across Hilbert space. This is often not something that can be probed by looking at correlation functions of local operators, which would be the preferred method for thinking about thermalization. While the bounds derived in this section do not mandate that the relaxation of local operators are controlled by ${\sf Dehn}_{g}(L)$, in Sec. IV we will explore an explicit example in which they are. Until then, we will focus on the more “global” timescales mentioned above. It is of course impossible to use ${\sf Dehn}_{g}(L)$—or any other quantity—to place upper bounds on thermalization timescales without making any additional assumptions about the details of ${\sf Dyn}_{G}$ (as without additional assumptions we could always choose ${\sf Dyn}_{G}$ to be evolution with a many body localized Hamiltonian, and the relevant timescales would all diverge).
@@ III.1 Circuit dynamics
where each $U_{j,g}$ acts on a length-$\ell_{R}$ block of sites and $G_{l}$ as before denotes the set of group elements expressible as words of length $\leq l$. Assuming the $U_{j,g}$ are drawn uniformly from the $|K_{g,\ell_{R}}|$-dimensional Haar ensemble,121212 Note that unlike in Ref. Singh et al. (2021), here we take the $U_{j,g}$ to be nontrivial in every $K_{g,\ell_{R}}$ sector, even the one-dimensional ones. This leads to a comparatively simpler expression for the Markov generator to appear below. performing the average gives
@@
as the projector onto the space of length-$\ell_{R}$ words $w$ with $\varphi(w)=g$, as well as—without loss of generality—taken $\mathcal{O}$ to factorize as $\mathcal{O}=\bigotimes_{j}\mathcal{O}_{j}$. Thus after a single step of the dynamics, all operators completely dephase and become diagonal in the computational basis, operators violating the dynamical constraint evaluate to zero under Haar averaging. We may thus focus on diagonal operators without loss of generality, which we will indicate using vector notation as $|\mathcal{O}(t)\rangle$. With this notation, (30) becomes
@@
$\mathcal{M}$ is a symmetric [^13] Strictly speaking $\mathcal{M}=\mathcal{M}^{T}$ only if we double each step of the evolution by including a reflection-related pattern of gates, viz. $\mathcal{M}\rightarrow\prod_{i=1}^{\ell_{R}}\mathcal{M}_{i}\prod_{j=1}^{\ell_{R}}\mathcal{M}_{\ell_{R}+1-j}$; we will tacitly assume this has always been done. doubly-stochastic matrix, with the smallest eigenvalue of 0 and the largest eigenvalue of 1. Due to the group constraint, $\mathcal{M}$ does not have a unique steady state following from that fact that the Markovian dynamics is reducible as
@@
A basic result Levin and Peres (2017) about $t_{\rm mix}(g)$ is that it is lower bounded by half the diameter of the state space that $\mathcal{M}_{g}$ acts on, simply because in order to mix, the system must at the minimum be able to traverse across most of state space. This gives the bound
@@
However, this bound is in fact too lose. Intuitively, this is because to saturate the bound, ${\sf Dyn}_{G}(t)$ would need to immediately find the optimal path between any two nodes in configuration space. Since $\mathcal{M}$ generates a random walk on state space, the dynamics will instead diffusively explore state space in a less efficient manner. On general grounds one might therefore expect a bound on $t_{\rm mix}(g)$ which is the square of the RHS above. This guess in fact turns out to be essentially correct Lyons and Peres (2017), with [^14] This follows from Prop. 13.7 of Lyons and Peres (2017) after using that in our case the equilibrium distribution of $\mathcal{M}_{g}$ is always uniform on $K_{g}$.
$t_{\rm mix}(g)\geq\frac{{\sf Dehn}_{g}^{2}(L)}{16\ln|K_{g,L}|}.$ (40)
Since $|K_{g,L}|$ is at most exponential in $L$, we thus have
@@
where $\lambda_{2}$ is the second-largest eigenvalue of $\mathcal{M}_{g}$’. This quantity admits a similar bound to $t_{\rm mix}(g)$, as one can show that Levin and Peres (2017)
@@ III.2 Hamiltonian dynamics
Thus far we have been talking about mixing times of Markov chains which encode the group dynamics. However, purely Hamiltonian dynamics does not mimic that of a Markov chain, and being reversible it possesses no direct analogues of mixing and relaxation times. Nevertheless, the Dehn function can still be used to place bounds on the timescales taken for time evolution to move wavefunctions around in Hilbert space. To be more quantitative, we will focus on the hitting time $t_{\rm hit}(g)$ of ${\sf Dyn}_{G}$, which we define as the minimum time needed for product states in $K_{g,L}$ to “reach” all other product states in $K_{g,L}$. Since however $\langle w^{\prime}|e^{-iHt}|w\rangle$ will generically be nonzero for all $|w\rangle,|w^{\prime}\rangle\in K_{g,L}$ as soon as $t>0$, we will need a slightly different definition of $t_{\rm hit}(g)$ as compared with the Markov chain case. To define $t_{\rm hit}(g)$ more precisely, we first define the hitting amplitude between two computational-basis product states $|w\rangle,|w^{\prime}\rangle\in K_{g,L}$ as
@@
Note that $h_{w,\cdot}(t;g)$ is a probability distribution on $K_{g,L}$, with $\sum_{w^{\prime}\in K_{g,L}}h_{ww^{\prime}}(t;g)=1$ for all $|w\rangle\in K_{g,L}$. We define $t_{\rm hit}(g,w,w^{\prime})$ between two words $w$ and $w^{\prime}$ as the minimum time for which $h_{ww^{\prime}}(t;g)$ reaches a fixed fraction of its infinite-time average defined by
@@ IV The Baumslag-Solitar group: exponentially slow hydrodynamics
The previous two sections have focused on developing parts of the general theory of group dynamics. In this section we take an in-depth look at a particular example which exhibits anomalously long relaxation times. Our example will come from a family of groups whose Dehn functions scale exponentially with $L$, yielding word problems with large time complexity. [^15] The spatial complexity in these examples—which will be discussed later in Sec. V.1—is however much smaller. When ${\sf Dehn}(L)\sim\exp(L)$, the hitting and mixing times discussed in the previous section are exponentially long. It is perhaps already rather surprising that a translation invariant local Markov process / unitary circuit can have mixing times that are this long, but our example is most interesting for another reason: it also possesses a global symmetry whose conserved charge takes ${\sf Dehn}(L)$ time to relax. This allows the slow time complexity of the word problem to be manifested in the expectation values of simple local operators, rather than being hidden in hitting times between different product states. Our example comes from a family of groups known in the math literature as the Baumslag-Solitar groups Baumslag (1969), which are parameterized by two integers $n,m\in\mathbb{N}$. Each group ${\sf BS}(n,m)$ in this family is generated by two elements $\mathtt{a}$ and $\mathtt{b}$, which obey the following relation:
@@
Models with ${\sf Dyn}_{{\sf BS}(n,m)}$ dynamics are thus most naturally realized in spin-2 systems with $\max(n,m)$-local dynamics. The simplest albeit interesting Baumslag-Solitar group is ${\sf BS}(1,2)$, and the rest of our discussion will focus on this case. For convenience, in what follows we will write ${\sf BS}(1,2)$ simply as ${\sf BS}$. The nontrivial relation in ${\sf BS}$ reads
@@ IV.1 A geometric perspective on group complexity
To understand how dynamics in ${\sf Dyn}_{\sf BS}$ works, we will find it helpful to introduce a geometric way of thinking about the group word problem. Given a general discrete group $G$, we will let ${\mathsf{CG}}_{G}$ denote the Cayley graph of $G$. Recall that ${\mathsf{CG}}_{G}$ is a graph with vertices labeled by elements of $G$ and edges labeled by generators of $G$ and their inverses, with two vertices $\mathtt{h},\mathtt{k}$ being connected by an edge $\mathtt{g}$ iff $\mathtt{h}=\mathtt{g}\mathtt{k}$. As simple examples, ${\mathsf{CG}}_{\mathbb{Z}_{N}}$ for $\mathbb{Z}_{N}=\langle\mathtt{g}\,|\,\mathtt{g}^{N}=\mathtt{e}\rangle$ is a length-$N$ closed cycle; ${\mathsf{CG}}_{\mathbb{Z}^{2}}$ for $\mathbb{Z}^{2}=\langle\mathtt{x},\,\mathtt{y}\,|\,\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}\rangle$ is a 2D square lattice; and ${\mathsf{CG}}_{\mathbb{Z}\ast\mathbb{Z}}$ for the free group $\mathbb{Z}\ast\mathbb{Z}=\langle\mathtt{x},\,\mathtt{y}\,|\,\rangle$ is a Bethe lattice with coordination number 4.161616While the exact structure of ${\mathsf{CG}}_{G}$ is different for different presentations of $G$, the large scale structure of ${\mathsf{CG}}_{G}$ (scalings of geodesic lengths, graph expansion, and so on) is presentation-independent (see e.g. Ref. Clay and Margalit (2017) for background). Therefore we will mostly refer to the Cayley graph of a group, rather than a particular presentation thereof.
@@
Figure 3: The group geometry of ${\sf BS}(1,2)$. The Cayley graph of ${\sf BS}(1,2)$ is constructed from an infinite number of two-dimensional sheets glued together in a tree-like fashion. a) A single sheet of the Cayley graph, which resembles a hyberbolic plane. Orange arrows denote multiplication by $a$ and purple arrows multiplication by $b$, with the boundary of each plaquette being the defining relation $aba^{-1}a^{-1}b^{-1}=e$. The diagonal purple lines denote edges which connect to different sheets; all diagonal edges connected to vertices at the same ”height” connect to the same sheet. b) How different sheets are glued together to form the full Cayley graph. The shaded yellow region denotes how a section of a single sheet is embedded in the full Cayley graph. The bold dashed path denotes an example of a word $w_{\mathrm{large}}(n)$ homotopic to the identity and with large area but small perimeter (in the figure, $n=3$).
@@
It is useful to realize that from any Cayley graph ${\mathsf{CG}}_{G}$, we can always construct a related 2-complex by associating oriented faces (or 2-cells) to each of ${\mathsf{CG}}_{G}$’s elementary closed loops; the 2-cells have the property that the product of generators around their boundary is a relation in $R$. For the just-mentioned examples, $\mathbb{Z}_{N}$ would have one $N$-sided face, $\mathbb{Z}^{2}$ would have a face for each plaquette of the square lattice (with the generators around the plaquette boundaries reading $\mathtt{x}\mathtt{y}\mathtt{x}^{-1}\mathtt{y}^{-1}$), and $\mathbb{Z}\ast\mathbb{Z}$ would have no faces at all (due to its lack of non-trivial relations). The 2-complex thus constructed is known as the Cayley 2-complex of $G$, and we will abuse notation by also referring to it as ${\mathsf{CG}}_{G}$. Any group word $w\in W(G)$ naturally defines an oriented path in ${\mathsf{CG}}_{G}$ obtained by starting at an (arbitrarily chosen) origin of ${\mathsf{CG}}_{G}$ and moving along edges based on the characters in $w$. The endpoint of this path on the Cayley graph is thus the group element $\varphi(w)$. Additionally, applying local relations in $R$ to $w$ deforms this path while keeping its endpoints fixed. This gives a geometric interpretation of the subspaces $K_{g,L}$:
@@
${\sf Dehn}(L)=\max_{\text{loops $w$ of length $L$}}{\sf Dehn}(w,\mathtt{e}^{L}).$ (67)

### 11 Apr 2024 v2

@@ III Slow thermalization and the time complexity of the word problem
From the discussion of the previous section, it is natural to expect that systems with ${\sf Dyn}_{G}$ dynamics will have thermalization times controlled by the time complexity of the word problem, as diagnosed by the Dehn functions ${\sf Dehn}_{g}(L)$ defined in (14). In this section we make this relationship precise by using the functions ${\sf Dehn}_{g}(L)$ to place lower bounds on various thermalization timescales. In the case of random unitary or classical stochastic dynamics, ${\sf Dehn}_{g}(L)$ will be used to bound relaxation and mixing times; for Hamiltonian dynamics it will appear in bounds for hitting times. A common feature these timescales have is that they quantify when ${\sf Dyn}_{G}$ is able to spread initial product states across Hilbert space. This is often not something that can be probed by looking at correlation functions of local operators, which would be the preferred method for thinking about thermalization. While the bounds derived in this section do not mandate that the relaxation of local operators also be controlled by ${\sf Dehn}_{g}(L)$, in Sec. IV we will explore an explicit example in which they are. Until then, we will focus on the more “global” diagnostics of relaxation and hitting times. We note in passing that it is impossible to use ${\sf Dehn}_{g}(L)$—or any other quantity—to place upper bounds on thermalization timescales without making any additional assumptions about the details of ${\sf Dyn}_{G}$ (as without additional assumptions we could always choose ${\sf Dyn}_{G}$ to be evolution with a many body localized Hamiltonian, and the relevant timescales would all diverge). Even in the case where ${\sf Dyn}_{G}$ is an appropriately constrained form of random unitary evolution, upper bounding the relaxation timescales requires techniques beyond those employed below, and constitutes an interesting direction for future work.
@@ III.1 Circuit dynamics
where each $U_{j,g}$ acts on a length-$\ell_{R}$ block of sites and $G_{\ell_{R}}$ as before denotes the set of group elements expressible as words of length $\leq\ell_{R}$. Assuming the $U_{j,g}$ are drawn uniformly from the $|K_{g,\ell_{R}}|$-dimensional Haar ensemble,131313 Note that unlike in Ref. [52], here we take the $U_{j,g}$ to be nontrivial in every $K_{g,\ell_{R}}$ sector, even the one-dimensional ones (where $U_{j,g}$ acts as multiplication by a random phase). This leads to a comparatively simpler expression for the Markov generator to appear below. performing the average gives
@@
as the projector onto the space of length-$\ell_{R}$ words $w$ with $\varphi(|w\rangle)=g$, as well as—without loss of generality—taken $\mathcal{O}$ to factorize as $\mathcal{O}=\bigotimes_{j}\mathcal{O}_{j}$. Thus after a single step of the dynamics, all operators completely dephase and become diagonal in the computational basis, operators violating the dynamical constraint evaluate to zero under Haar averaging. We may thus focus on diagonal operators without loss of generality, which we will indicate using vector notation as $|\mathcal{O}(t)\rangle$. With this notation, (32) becomes
@@
$\mathcal{M}$ is a symmetric [^14] Strictly speaking $\mathcal{M}=\mathcal{M}^{T}$ only if we double each step of the evolution by including a reflection-related pattern of gates, namely $\mathcal{M}\rightarrow\prod_{i=1}^{\ell_{R}}\mathcal{M}_{i}\prod_{j=1}^{\ell_{R}}\mathcal{M}_{\ell_{R}+1-j}$; we will tacitly assume this has always been done. doubly-stochastic matrix, with the smallest eigenvalue of 0 and the largest eigenvalue of 1. Due to the group constraint, $\mathcal{M}$ does not have a unique steady state following from that fact that the Markovian dynamics is reducible as
@@
A basic result [53] about $t_{\rm mix}(g)$ is that it is lower bounded by half the diameter of the state space that $\mathcal{M}_{g}$ acts on, simply because in order to mix, the system must at the minimum be able to traverse across most of state space. This gives the bound
@@
However, this bound is in fact too lose. Intuitively, this is because to saturate the bound, ${\sf Dyn}_{G}(t)$ would need to immediately find the optimal path between any two nodes in configuration space. Since $\mathcal{M}$ generates a random walk on state space, the dynamics will instead diffusively explore state space in a less efficient manner. On general grounds one might therefore expect a bound on $t_{\rm mix}(g)$ which is the square of the RHS above. This guess in fact turns out to be essentially correct, with
$t_{\rm mix}(g)\geq\frac{{\sf Dehn}_{g}^{2}(L)}{16\ln|K_{g,L}|},$ (42)
which follows from Prop. 13.7 of [54] after using that the equilibrium distribution of $\mathcal{M}_{g}$ is always uniform on $K_{g}$ by virtue of $\mathcal{M}_{g}$ being doubly stochastic. Since $|K_{g,L}|$ is at most exponential in $L$, we thus have
@@
where $\lambda_{2}$ is the second-largest eigenvalue of $\mathcal{M}_{g}$. This quantity admits a similar bound to $t_{\rm mix}(g)$, as one can show that [53]
@@ III.2 Hamiltonian dynamics
We have thus far discussed mixing times of the Markov chains which arise from random unitary dynamics. However, purely Hamiltonian dynamics does not mimic that of a Markov chain, and being reversible it possesses no direct analogues of mixing and relaxation times. Nevertheless, the Dehn function can still be used to place bounds on the timescales taken for time evolution to spread wavefunctions across in Hilbert space. To be more quantitative, we will focus on the hitting time $t_{\rm hit}(g)$ of ${\sf Dyn}_{G}$, which we define as the minimum time needed for product states in $K_{g,L}$ to “reach” all other product states in $K_{g,L}$. Since $\langle w^{\prime}|e^{-iHt}|w\rangle$ will generically be nonzero for all $|w\rangle,|w^{\prime}\rangle\in K_{g,L}$ as soon as $t>0$, we will need a slightly different definition of $t_{\rm hit}(g)$ as compared with the Markov chain case. To define $t_{\rm hit}(g)$ more precisely, we first define the hitting amplitude between two computational-basis product states $|w\rangle,|w^{\prime}\rangle\in K_{g,L}$ as
@@
Note that $h_{w,\cdot}(t;g)$ is a probability distribution on $K_{g,L}$, with $\sum_{w^{\prime}\in K_{g,L}}h_{ww^{\prime}}(t;g)=1$ for all $|w\rangle\in K_{g,L}$. We define $t_{\rm hit}(g,w,w^{\prime})$ between two words $w$ and $w^{\prime}$ as the minimum time for which $h_{ww^{\prime}}(t;g)$ reaches a fixed fraction of its infinite-time average $\mkern 1.5mu\overline{\mkern-1.5muh_{ww^{\prime}}\mkern-1.5mu}\mkern 1.5mu(g)$, defined by
@@ IV The Baumslag-Solitar group: exponentially slow hydrodynamics
The previous two sections have focused on developing parts of the general theory of semigroup dynamics. In this section we take an in-depth look at a particular example which exhibits anomalously long relaxation times. Our example will come from a family of groups whose Dehn functions scale exponentially with $L$, yielding word problems with large time complexity. [^15] The spatial complexity of these examples—which will be discussed later in Sec. V.1—is however trivial. When ${\sf Dehn}(L)\sim\exp(L)$, the hitting and mixing times discussed in the previous section are exponentially long. It is perhaps already rather surprising that a translation invariant local Markov process / unitary circuit can have mixing times that are this long, but our example is most interesting for another reason: it also possesses a global symmetry whose conserved charge takes ${\sf Dehn}(L)$ time to relax. This allows the slow time complexity of the word problem to be manifested in the expectation values of simple local operators, rather than being hidden in hitting times between different product states. Our example comes from a family of groups known in the math literature as the Baumslag-Solitar groups [55], which are parameterized by two integers $n,m\in\mathbb{N}$. Each group ${\sf BS}(n,m)$ in this family is generated by two elements $\mathtt{a}$ and $\mathtt{b}$ obeying a single nontrivial relation, with the group presentation
@@
Models with ${\sf Dyn}_{{\sf BS}(n,m)}$ dynamics are thus most naturally realized in spin-2 systems with $\max(n,m)$-local dynamics and onsite Hilbert space
$\mathcal{H}_{\rm loc}={\rm span}\{|\mathtt{a}\rangle,\,|\mathtt{a}^{-1}\rangle,\,|\mathtt{b}\rangle,\,|\mathtt{b}^{-1}\rangle,\,|\mathtt{e}\rangle\}.$ (62)
The simplest Baumslag-Solitar group which is interesting for our purposes is ${\sf BS}(1,2)$, and the rest of our discussion will focus on this case. For convenience, in what follows we will write ${\sf BS}(1,2)$ simply as ${\sf BS}$. The nontrivial relation in ${\sf BS}$ reads
@@ IV.1 A geometric perspective on group complexity
To understand how dynamics in ${\sf Dyn}_{\sf BS}$ works, we will find it helpful to introduce a geometric way of thinking about the group word problem. [^16] This picture is most useful for groups, rather than semigroups; see Sec. VI and App. A for discussions on this point. Given a general discrete group $G$, we will let ${\mathsf{CG}}_{G}$ denote the Cayley graph of $G$. Recall that ${\mathsf{CG}}_{G}$ is a graph with vertices labeled by elements of $G$ and edges labeled by generators of $G$ and their inverses, with two vertices $\mathtt{h},\mathtt{k}$ being connected by an edge $\mathtt{g}$ iff $\mathtt{h}=\mathtt{g}\mathtt{k}$. As simple examples, ${\mathsf{CG}}_{\mathbb{Z}_{N}}$ for $\mathbb{Z}_{N}=\langle\mathtt{g}\,|\,\mathtt{g}^{N}=\mathtt{e}\rangle$ is a length-$N$ closed cycle; ${\mathsf{CG}}_{\mathbb{Z}^{2}}$ for $\mathbb{Z}^{2}=\langle\mathtt{x},\,\mathtt{y}\,|\,\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}\rangle$ is a 2D square lattice; and ${\mathsf{CG}}_{\mathbb{Z}\ast\mathbb{Z}}$ for the free group $\mathbb{Z}\ast\mathbb{Z}=\langle\mathtt{x},\,\mathtt{y}\,|\,\rangle$ is a Bethe lattice with coordination number 4.171717While the exact structure of ${\mathsf{CG}}_{G}$ is different for different presentations of $G$, the large scale structure of ${\mathsf{CG}}_{G}$ (scalings of geodesic lengths, graph expansion, and so on) is presentation-independent (see e.g. Ref. [41] for background). Therefore we will mostly refer to the Cayley graph of a group, rather than a particular presentation thereof.
@@
Figure 3: The group geometry of ${\sf BS}(1,2)$. The Cayley graph of ${\sf BS}(1,2)$ is constructed from an infinite number of two-dimensional sheets glued together in a tree-like fashion. a) A single sheet of the Cayley graph, which resembles a hyberbolic plane. Orange arrows denote multiplication by $\mathtt{a}$ and purple arrows multiplication by $\mathtt{b}$, with the boundary of each plaquette being the defining relation $\mathtt{a}\mathtt{b}\mathtt{a}^{-1}\mathtt{a}^{-1}\mathtt{b}^{-1}=e$. The diagonal purple lines denote edges which connect to different sheets; all diagonal edges connected to vertices at the same ”height” connect to the same sheet. b) How different sheets are glued together to form the full Cayley graph. The shaded yellow region denotes how a section of a single sheet is embedded in the full Cayley graph. The bold dashed path denotes the path traced out by a “large” word $w_{\mathrm{large}}(n)=\mathtt{a}\mathtt{b}^{-n}\mathtt{a}^{-1}\mathtt{b}^{n}\mathtt{a}^{-1}\mathtt{b}^{-n}\mathtt{a}\mathtt{b}^{n}$, which is homotopic to the identity ($\varphi(|w_{\rm large}\rangle)=e$) and possesses an area scaling exponentially with its length. In the figure $n=3$, and the path on the Cayley graph is ordered by points 1 to 8, according to the 8 different components of $w_{\rm large}$ (read from right to left). Thus the point 2 corresponds to the word $\mathtt{b}^{3}$, the point 5 corresponds to $\mathtt{a}^{-1}\mathtt{b}^{-3}\mathtt{a}\mathtt{b}^{3}$, and so on.
@@
It is useful to realize that from any Cayley graph ${\mathsf{CG}}_{G}$, we can always construct a related 2-complex by associating oriented faces (or 2-cells) to each of ${\mathsf{CG}}_{G}$’s elementary closed loops; the 2-cells have the property that the product of generators around their boundary is a relation in $R$. For the just-mentioned examples, $\mathbb{Z}_{N}$ would have one $N$-sided face, $\mathbb{Z}^{2}$ would have a face for each plaquette of the square lattice (with the generators around the plaquette boundaries reading $\mathtt{x}\mathtt{y}\mathtt{x}^{-1}\mathtt{y}^{-1}$), and the free product $\mathbb{Z}\ast\mathbb{Z}$ would have no faces at all (due to its lack of non-trivial relations). The 2-complex thus constructed is known as the Cayley 2-complex of $G$, and we will abuse notation by also referring to it as ${\mathsf{CG}}_{G}$. Any group word $w\in W(G)$ naturally defines an oriented path in ${\mathsf{CG}}_{G}$ obtained by starting at an (arbitrarily chosen) origin of ${\mathsf{CG}}_{G}$ and moving along edges based on the characters in $w$. The endpoint of this path on the Cayley graph is thus the group element $\varphi(|w\rangle)$. Additionally, applying local relations in $R$ to $w$ deforms this path while keeping its endpoints fixed. This gives a geometric interpretation of the subspaces $K_{g,L}$:
@@
${\sf Dehn}(L)=\max_{\text{loops {\hbox{w}} of length {\hbox{L}}}}{\sf Dehn}(w,\mathtt{e}^{L}).$ (70)